OpenStudy (anonymous):

4e - 8= 32? It says to find the variable?

OpenStudy (anonymous):

the variable is e e = 10

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

4e - 8 + 8 = 32 + 8 4e = 40 4e/4 = 40/4 e = 10

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

A + 8 = -32 it says to find the variable?

OpenStudy (anonymous):

How did you get A?

OpenStudy (anonymous):

idk it says A+18 = -32

OpenStudy (anonymous):

ur confusing me, you said "4e - 8= 32? It says to find the variable? "

OpenStudy (anonymous):

No its a different problem

OpenStudy (anonymous):

ohhh okay lol umm so the variable is A a + 18 - 18 = -32 - 18 a = -40

OpenStudy (anonymous):

Are you sure? because i dont have a choice of 40 all i have is a choice of 36

OpenStudy (anonymous):

- amd -1 and 9

OpenStudy (anonymous):

ur questions A + 18 = -32 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so: A + 18 - 18 = -32 - 18 A = -50 sorry not -40 idunno anythin but these r my best guesses

OpenStudy (aroub):

Yep, -50 :)

OpenStudy (anonymous):

-50